2. Leak Inc. forecasts the free cash flows (in millions) shown below. If the weighted average cost of capital is 11% and FCF is expected to grow at a rate of 5% after Year 2, what is the Year 0 value of operations, in millions? Assume that the ROIC is expected to remain constant in Year 2 and beyond (and do not make any half-year adjustments).
Year: 1 2
Free cash flow: -$50 $100

a. $1,456
b. $1,529
c. $1,606
d. $1,686
e. $1,770
1
Expert's answer
2012-02-16T09:31:01-0500
Terminal Value in year 2 = CF3/(Ke-g) where
CF3 = cash flow in year 3 = 100X1.05 = 105
Ke = required return = 11%
g = growth rate = 5%
Terminal value in year 2 = 105/(11%-5%) = 1,770
e. $1,770